# Resilient and Freshness-Aware Scheduling for Industrial Multi-Hop IAB Networks: A Packet Duplication Approach

### v1

@@ I Introduction
To enhance transmission reliability, the 3rd Generation Partnership Project (3GPP) introduced the Packet Duplication (PD) strategy, which transmits critical data packets through disjoint paths [3]. Although PD significantly enhances the resilience against single-path failures, it fundamentally violates the goal of resource efficiency and inevitably exacerbates congestion and queuing delays [4]. Therefore, there exists a tricky trade-off between the traffic load in the network and the freshness of the data.
@@
In single-hop networks, the AoI-oriented problems has been well studied. In recent years, the research focus of researchers has shifted to multi-hop networks, especially in IAB networks. Existing literature has extensively explored relay-aided AoI minimization strategies [5, 6], multi-path scheduling designed for high-capacity transmission [7], and energy-efficient resource allocation in IIoT [8]. However, the specific intersection of PD strategy and freshness-aware scheduling in IAB networks remains largely unexplored. Unlike the aforementioned studies that consider throughput capacity or energy savings, deploying PD for reliability introduces unique challenges. The redundant traffic imposes severe pressure on resource constrained multi-hop networks. Therefore, a critical gap remains in addressing the specific trade-off between reliability guarantees and the reduction in freshness caused by congestion.
@@ II-A Mathematical Modeling
A typical scenario of a resilient industrial IAB network is shown in Fig.1, which contains an IAB donor, $\mathit{V}$ IAB nodes, and $\mathit{U}$ User Equipments (UEs), e.g. AGVs, uniformly deployed within an industrial factory. The network topology is modeled as a directed graph $\mathcal{G}=\{\mathcal{N},\mathcal{L}\}$, where $\mathcal{N}=\mathcal{V}\cup\mathcal{U}$, $\mathcal{V}=\{0,...,\mathit{V}\}$ represents the set of IAB nodes with the IAB donor indexed as 0, $\mathcal{U}=\{1,...,\mathit{U}\}$ represents the UE set, and $\mathcal{L}=\{l_{i,j}|i,j\in\mathcal{N}\}$ stands for the set of directed wireless links, where $l_{i,j}$} and $l_{j,i}$} denoting distinct unidirectional links.
@@ II-A4 Scheduling Constraints
Furthermore, leveraging the high directionality and severe path loss of mmWave communications, we assume interference is negligible for nodes separated by a distance greater than $R$, as mmWave networks typically operate in a noise-limited regime due to blockage and narrow beamforming [10]. Within this range, we enforce an interference avoidance scheduling constraints. By integrating this with the half-duplex constraint of IAB node, where simultaneous transmission and reception are prohibited. We construct a conflict set $\mathcal{I}_{i,j}$ for each link $l_{i,j}$. This set encompasses all potential links that either are located within the interference range $R$ or share a common node with link $l_{i,j}$. To guarantee conflict-free transmission, at most one link within the conflict set and the link $l_{i,j}$ itself can be activated simultaneously at any time slot $t$, which is constrained by:
@@ II-C Problem Reformulation
$\begin{aligned} (\textbf{P2})\max_{\mathbf{X}}\sum_{l_{i,j}\in\mathcal{L}}\sum_{f\in\mathcal{F}}&x_{i,j,f}(t)\cdot\\ &\left[\gamma\left(Q_{i}^{f}(t)-Q_{j}^{f}(t)\right)+\Phi_{i,j,f}^{\text{AoI}}(t)\right]\end{aligned}$ (12)
s.t.  routing constraints:  Eq.(7)(8);
    scheduling constraints:  Eq.(9)(10),
where $\Phi_{i,j,f}^{\text{AoI}}(t)$ represents the expected AoI reduction offered by scheduling flow $f$ on link $l_{i,j}$. To encourage multi-hop packet forwarding at intermediate nodes, we define the gain term as a global urgency indicator:
$\Phi_{i,j,f}^{\text{AoI}}(t)\triangleq A_{d_{f},f}(t),\forall l_{i,j}\in\mathcal{L}.$ (13)
@@ III Proposed Low-Complexity Solution
For each time slot $t$, the algorithm initializes an empty candidate activation list $\mathcal{K}$. The scheduler evaluates the dynamic priority for every link $l_{i,j}$ associated with flow $f$ in the route path set $P_{f}$. Guided by the objective derived in Problem P2, the scheduling weight $W_{i,j,f}(t)$ is formulated as:
$W_{i,j,f}(t)=$ $\gamma\cdot\left(Q_{i}^{f}(t)-Q_{j}^{f}(t)\right)+A_{d_{f},f}(t),$ (14)
$\forall f\in\mathcal{F},\forall l_{i,j}\in P_{f},1\leq t\leq T.$
@@ IV-A Simulation Setup
We have constructed an industrial mmWave IAB network comprising a central IAB donor and 24 IAB nodes. These nodes are deployed in a regular grid pattern over an 800 m $\times$ 800 m factory area with an inter-node spacing of 200 meters. The interference range $R$ is defined as 100 m. In our simulations, the control parameter $\gamma$ is fixed at 0.5. This value was empirically determined to offer the best AoI performance subject to the zero-overflow constraint, as lower values triggered instability while higher values induced unnecessary waiting delays.
@@
The performance of the proposed RFAS algorithm is compared against three baselines. Firstly, there was a single-path scheme without considering PD. Additionally, there were two algorithms based on the MAX-weight scheduling concept, including the Queue-Aware Scheduling (QAS) scheme [13], which prioritizes queue stability and load balancing , and the Freshness-Aware Scheduling (FAS) scheme [9], which prioritizes AoI.
@@ IV-B2 Micro-Dynamics and Stability Trade-off
(a) Micro-Dynamics: AoI
(b) Micro-Dynamics: Stability
Figure 4: Micro-Dynamics Performance evaluation showing QAS, FAS, and RFAS. (a) AoI; (b) Queue Stability.
We introduce a strict buffer capacity constraint of $\Gamma_{n}=8$ packets, representing finite buffer size in low-cost IAB nodes. We analyze the transient response to a sudden traffic burst injected at $t=20$, as depicted in Fig.4a and Fig.4b. Since FAS only focuses on reducing the negative impacts of AoI, it is unable to promptly clear the backlog queue, as shown in Figure 4b. Therefore, although FAS has a theoretically novel advantage, it is not applicable to resource-constrained industrial multi-hop IAB network scenarios. In contrast, the proposed RFAS can effectively enforce strict constraints. Once the queue length approaches the upper limit, the system triggers forced transmission to alleviate congestion. This mechanism successfully limits the worst-case queue length, achieving a load balancing comparable to the QAS. Crucially, Fig.4a indicates that this strict stability enforcement incurs a lower cost. RFAS can ensure physical stability while reducing AoI.
@@ IV-B3 Scalability and Industrial Robustness
(a) Scalability: AoI
(b) Scalability: Imbalance
Figure 5: Scalability Performance evaluation showing QAS, FAS, and RFAS. (a) AoI; (b) Imbalance.
We verified the scalability of the algorithm and its performance in handling actual industrial traffic models. The results of scalability are shown in Fig.5a and Fig.5b. When the number of UEs increased from 4 to 14, the AoI of QAS significantly surged by approx $300\%$, while FAS maintained the lowest AoI, but it had a serious load imbalance problem, and the standard deviation of the queue length peaked at 3.5. In contrast, RFAS maintained a balanced performance. It effectively suppressed queue congestion while yielding an AoI within $10\%$ of the FAS lower bound. It is important to note that FAS achieved its theoretical minimum AoI only by disregarding the hard buffer constraints, which inevitably leads to physical overflows. From the standard deviation of approximately 2.8 in Fig.5b, RFAS improves load balancing by $25\%$ compared to FAS, it can be seen that compared to methods that solely focus on freshness, RFAS has significant improvements in load balancing.

### v2

@@ I Introduction
To enhance transmission reliability, the 3rd Generation Partnership Project (3GPP) introduced the Packet Duplication (PD) strategy, which transmits critical data packets through disjoint paths [3]. Although PD significantly enhances the resilience against single-path failures, it fundamentally violates the goal of resource efficiency and inevitably exacerbates congestion and queuing delays [4, 5]. Therefore, there exists a tricky trade-off between the traffic load in the network and the freshness of the data.
@@
In single-hop networks, the AoI-oriented problems has been well studied. In recent years, the research focus of researchers has shifted to multi-hop networks, especially in IAB networks. Existing literature has extensively explored relay-aided AoI minimization strategies [6, 7, 8], multi-path scheduling designed for high-capacity transmission [9, 10], and energy-efficient resource allocation in IIoT [11, 12]. However, the specific intersection of PD strategy and freshness-aware scheduling in IAB networks remains largely unexplored. Unlike the aforementioned studies that consider throughput capacity or energy savings, deploying PD for reliability introduces unique challenges. The redundant traffic imposes severe pressure on resource constrained multi-hop networks. Therefore, a critical gap remains in addressing the specific trade-off between reliability guarantees and the reduction in freshness caused by congestion.
@@ II-A Mathematical Modeling
A typical scenario of a resilient industrial IAB network is shown in Fig.1, which contains an IAB donor, $\mathit{V}$ IAB nodes, and $\mathit{U}$ User Equipments (UEs), e.g., AGVs, uniformly deployed within an industrial factory. The network topology is modeled as a directed graph $\mathcal{G}=\{\mathcal{N},\mathcal{L}\}$, where $\mathcal{N}=\mathcal{V}\cup\mathcal{U}$, $\mathcal{V}=\{0,...,\mathit{V}\}$ represents the set of IAB nodes with the IAB donor indexed as 0, $\mathcal{U}=\{1,...,\mathit{U}\}$ represents the UE set, and $\mathcal{L}=\{l_{i,j}|i,j\in\mathcal{N}\}$ stands for the set of directed wireless links, where $l_{i,j}$} and $l_{j,i}$} denoting distinct unidirectional links.
@@ II-A4 Scheduling Constraints
Furthermore, we assume that the scheduler obtains real-time blockage status through periodic beam training and rapid link-layer feedback. Leveraging the high directionality and severe path loss of mmWave communications, we assume interference is negligible for nodes separated by a distance greater than $R$. While dense deployments may exhibit complex interference, this model effectively captures the dominant conflict-set dynamics, as mmWave networks typically operate in a noise-limited regime due to blockage and narrow beamforming [14]. Within this range, we enforce an interference avoidance scheduling constraints. By integrating this with the half-duplex constraint of IAB node, where simultaneous transmission and reception are prohibited. We construct a conflict set $\mathcal{I}_{i,j}$ for each link $l_{i,j}$. This set encompasses all potential links that either are located within the interference range $R$ or share a common node with link $l_{i,j}$. To guarantee conflict-free transmission, at most one link within the conflict set and the link $l_{i,j}$ itself can be activated simultaneously at any time slot $t$, which is constrained by:
@@ III Proposed Low-Complexity Solution
For each time slot $t$, the algorithm initializes an empty candidate activation list $\mathcal{K}$. The scheduler evaluates the dynamic priority for every link $l_{i,j}$ associated with flow $f$ in the route path set $P_{f}$.
@@
Although we adopted an offline configuration of shortest paths to reduce complexity, in each time slot of the online phase, the algorithm can dynamically select the optimal available path based on the link status. In extreme cases, if all candidate paths are blocked, the queue stability constraint can ensure that data packets are buffered until a feasible path is restored, thereby maintaining the resilience of the system.
@@ IV-A Simulation Setup
We have constructed an industrial mmWave IAB network comprising a central IAB donor and 24 IAB nodes. These nodes are deployed in a regular grid pattern over an 800 m $\times$ 800 m factory area with an inter-node spacing of 200 meters. The interference range $R$ is defined as 100 m. We evaluated the sensitivity of the control parameter $\gamma$, which reflects the trade-off between freshness and stability. Specifically, a larger $\gamma$ will prioritize the AoI of data packets, but it will increase the occupancy rate of the buffer; while a smaller $\gamma$ can improve the stability of the queue, but it will lead to higher AoI; therefore, through experiments, we determined that $\gamma=0.5$ is the optimal ideal balance point, which can minimize AoI while avoiding buffer overflow or system instability.
@@
The performance of the proposed RFAS algorithm is compared against three baselines. Firstly, there was a single-path scheme without considering PD. Additionally, there were two algorithms based on the MAX-weight scheduling concept, including the Queue-Aware Scheduling (QAS) scheme [17], which prioritizes queue stability and load balancing , and the Freshness-Aware Scheduling (FAS) scheme [13], which prioritizes AoI. By comparing RFAS with these baselines, we can effectively capture the trade-off between reliability and freshness scheduling, thereby conducting a comprehensive assessment of the performance of the proposed algorithm.
@@ IV-B2 Micro-Dynamics and Stability Trade-off
We introduce a strict buffer capacity constraint of $\Gamma_{n}=8$ packets, representing finite buffer size in low-cost IAB nodes. We analyze the transient response to a sudden traffic burst injected at $t=20$. Since FAS only focuses on reducing the negative impacts of AoI, it is unable to promptly clear the backlog queue, as shown in . Therefore, although FAS has a theoretically novel advantage, it is not applicable to resource-constrained industrial multi-hop IAB network scenarios. In contrast, the proposed RFAS can effectively enforce strict constraints. Once the queue length approaches the upper limit, the system triggers forced transmission to alleviate congestion. This mechanism successfully limits the worst-case queue length, achieving a load balancing comparable to the QAS. Crucially, indicates that this strict stability enforcement incurs a lower cost. RFAS can ensure physical stability while reducing AoI.
@@ IV-B3 Scalability and Industrial Robustness
We verified the scalability of the algorithm and its performance in handling actual industrial traffic models. The results of scalability. When the number of UEs increased from 4 to 14, the AoI of QAS significantly surged by approx $300\%$, while FAS maintained the lowest AoI, but it had a serious load imbalance problem, and the standard deviation of the queue length peaked at 3.5. In contrast, RFAS maintained a balanced performance. It effectively suppressed queue congestion while yielding an AoI within $10\%$ of the FAS lower bound. It is important to note that FAS achieved its theoretical minimum AoI only by disregarding the hard buffer constraints, which inevitably leads to physical overflows. From the standard deviation of approximately 2.8, RFAS improves load balancing by $25\%$ compared to FAS, it can be seen that compared to methods that solely focus on freshness, RFAS has significant improvements in load balancing.